OpenStudy (anonymous):

calculate the area of the rectangle for each value of n. n=1 meter n=5 meters n=10 meters n=0.5 meter n=(5+m) meters sketch a rectangle

OpenStudy (anonymous):

U mean square?

OpenStudy (anonymous):

rectangle

OpenStudy (anonymous):



OpenStudy (anonymous):

AREA of a RECTANGLE =length * breadth

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

the meters and numbers that i put it should be calculated :D

OpenStudy (anonymous):

Now, when n=1 Area of rectangle = n*6=1*6=6 sq.meters

OpenStudy (anonymous):

Similarly, can u do the rest

OpenStudy (anonymous):

the rectangle i draw is to write a formula for the area of the rectangle in terms of n :)

OpenStudy (anonymous):

Its 6n

OpenStudy (anonymous):

oww ok